You are characterizing the binding of a ligand to a protein. Knowing that the rate of association between the protein and ligand is diffusion controlled and that the KD is expected to be around 100 µM,

a) What do you expect that half-life of the complex to be?

b) What method would you choose to measure the KD? Briefly describe why the method is appropriate
Answer #1

To characterize the binding of ligand to protein, measurement of association rate constant (K1) and dissociation rate constant (K-1) needed to estimate the equilibrium dissociation constant (KD)

KD = (K-1/K1)        ---------------------(1)

a) Half life of the protein - ligand complex (T1/2) = 0.693/K-1   ----------------------------(2)

Given, KD = 100µM,

if diffusion controlled association K1≈107 M-1 sec-1 ------------------(3)

substitute (3) in (1)

K-1 =    100 x 10 -6 M x 107 M-1 sec-1

K-1= 1000 sec-1           ---------------------(4)

substitute (4) in (2)

Half life of the complex expected to be 7 msec.

(b) Brownian dynamics simulations method used to calculate the diffusion-controlled rate constant kD. Because this method easy to comphrend for different protein shapes and various interaction potentials. In compared to analytical theories, have structural assumption are simple for evaluate transient complexes.